# Privacy Reasoning in Ambiguous Contexts

### v1

@@ 5.1 Enabling context disambiguation across datasets
Our context expansions are designed to simulate real-life clarifications a personal agent could obtain from a user, which can be label-independent (e.g., the data sender John is a mid-level manager at a startup) or label-dependent (i.e., the data subject Emily gave John her consent to share her information). In label-dependent expansions, we generate contexts designed to align with the example’s appropriateness label. Specifically, we generate inappropriate contexts (denoted $\mathcal{N^{-}}$) for $\mathcal{S^{-}}$ examples to reinforce their inappropriateness and appropriate contexts (denoted $\mathcal{N^{+}}$) for $\mathcal{S^{+}}$ examples to reinforce their appropriateness. Crucially, these label-dependent expansions, which deliberately reinforce the example’s initial appropriateness label, are distinct from user clarifications in authentic user-agent interactions. While real-word user clarifications typically aim to bolster agent decision confidence and can shift the perceived appropriateness in either direction, our expansions are not intended to mirror such interactions. Instead, they are specifically constructed to systematically evaluate how LLM privacy decision-making responds to additional contextual information generated via various strategies.
@@ 5.2 Reasoning-guided expansion
The reasoning that models provide alongside their appropriateness judgments reveals that models frequently make assumptions when context is under-specified. In this section we aim to uncover these implicit assumptions and subsequently generate context expansions that obviate model reliance on them. As we show in Section 6, this enables models to become significantly more effective in their privacy assessments.
@@
Suitability of communication channel (Suitability) Whether sending the data over this specific communication channel is okay $49$ 00
@@
Alignment with norms (Norms) Whether it is a standard practice to share this data in this context $29$ $7$
Purpose (Purpose) Whether there is an intended goal or specific reason for sharing or utilizing the information $16$ $7$
Recipient authorization (Recipient Auth) Whether the intended recipient is permitted to access or receive the information $15$ 00
Established practices (Practices) Whether sharing this information follows a previous approach or agreement between the parties $11$ $4$
Sender authorization (Sender Auth) Whether the sender is permitted to share or transmit the information $5$ $2$
Safety guidelines (Safety) Whether this information violates the content moderation policies of the platform on which it’s being shared $3$ 00
@@
Building on extensive research into LLM reasoning geva2021aristotle ; sprague2023musr ; yuchen2025zebralogic ; jacovi2024chain and factors influencing their outputs, we seek to understand appropriateness decisions of LLMs by prompting them to provide concise reasoning alongside each yes/no decision (our prompts can be found in Appendix E). Although elicited reasoning may not perfectly represent internal processes for models, it offers valuable insights into their decision-making chen2025reasoning . This reasoning elicitation step yielded performance comparable to that of models not providing explicit reasoning (P=83.5%, R=74.4%, and $F_{1}$=78.7for PrivacyLens+, and P=97.6%, R=75.9%, and $F_{1}$=85.4for ConfAIde+).
To understand model reasoning and identify implicit assumptions, two researchers analyzed 40 outputs for PrivacyLens+ (10 per category: TPs, TNs, FPs, FNs) and 21 outputs for ConfAIde+ (7 each for TPs, TNs, FNs; excluding FPs due to insufficient samples). They examined the prompts and reasoning for those examples—blind to ground-truth labels and LLM judgments—to identify assumptions and develop the nine codes shown in Table 2. Subsequently, using this established codebook, the same two researchers independently coded a larger, stratified random sample of 120 PrivacyLens+ examples (30 per category) and 50 ConfAIde+ examples (15 TPs, TNs, FNs, and 5 FPs). Disagreement was resolved through discussion to achieve consensus cascio2019team . Table 2 presents the resulting code occurrences across datasets.
@@
To test if codes are more effective at reducing implicit reasoning assumptions, we prompt the LLM to generate label-dependent contexts based on each code (denoted as $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$) for all $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples respectively in both datasets. Our instruction specifically asks the LLM to first identify the most suitable field to expand based on the code and its definition, and then to expand the selected field following this code to make the $\mathcal{S^{-}}$ more inappropriate, or $\mathcal{S^{+}}$ more appropriate. Similarly to Shao et al. salt-nlp2024privacylens , we explicitly ask the LLM to add descriptive contexts and avoid using evaluative words like “sensitive” and “non-sensitive” that would introduce reasoning shortcuts in the expansions. As we did for our label-dependent expansion of each field, for ConfAIde+ we again need to further integrate $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$ into $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$. In total, we generate an additional $9\times 493=$ 4,437 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples for PrivacyLens+, and $9\times 270=$ 2,430 pairs for ConfAIde+. We provided detailed descriptions of our generation procedure in Appendix D.1.3 and  D.2.4. In total, we generate 19,720 examples for PrivacyLens+, and 12,960 examples for ConfAIde+. We intend to make our datasets public upon paper acceptance.
@@ 6 Results
To measure the benefits of reasoning-guided contextual disambiguation we ask the following questions: (1) how reasoning-guided expansion compare to other expansion strategies, (2) whether reasoning-guided expansion improves privacy assessments and prompt sensitivity across models and prompt variants (3) what the sources of context ambiguity and drivers of performance gains are in reasoning-guided expansion, and (4) how well the reasoning-guided expansion codes generalize.
@@
Figure 4 compares the performance of privacy assessments on the reasoning-guided expansion with that on the original scenarios, and the label-independent and label-dependent expansions. For each expansion, we report the average $F_{1}$ across all fields/codes for that expansion and the $F_{1}$ for the top-performing field/code. Looking at the average expansion performance, we observe that the label-independent expansion provides little utility in improving performance over the no expansion baseline. In contrast, the label-dependent expansion shows consistent improvements across models, especially for PrivacyLens+. This is likely due to the expansion being based on contextual integrity: forcing context disambiguation to operate within the contextual integrity parameters provides clear benefits in improving appropriateness judgments. Nevertheless, the reasoning-guided expansion outperforms the others for both datasets, which further reinforces the benefits of basing context disambiguation on the codes that we introduced in Section 5.2.
@@
Figure 4 also highlights the performance of the top-performing field / code for each expansion, revealing a high degree of consistency in the fields / codes chosen by each model. On PrivacyLens+ dataset, the agreement is unanimous, with all three models having identical top-performing fields / codes for each respective strategy. This pattern of consensus is also observed for ConfAIde+, where there is strong agreement for both the label-dependent and reasoning-guided expansions.
@@
Analyzing the performance of the top-performing fields and codes clearly highlights the utility of reasoning-based expansion: Camber delivers substantial performance increase and reduction in prompt sensitivity that is also consistent across models, compared to the other expansion strategies. In addition, the expansion codes that we develop using PrivacyLens+ are also able to increase performance on the ConfAIde+ dataset, highlighting their generalization capability. A detailed breakdown and discussion of the performance across individual fields and codes is available in Appendix B.1B.2 and B.3.
Figure 5: $F_{1}$ scores across models and prompt variants when expanding with the four most prevalent codes. These codes show consistent performance improvements and prompt sensitivity reduction for all three models. Error bars show 95% confidence intervals generated by bootstrapping the experiment results 1,000 times with replacement.
Benefits of reasoning-guided expansions.
To further analyze the utility of adding reasoning-guided context, we compare the performance of Camber across the top 4 most prevalent codes in Table 2 for each dataset (Figure 5). We observe substantial performance gains and prompt sensitivity reductions compared to the baselines across all three models. These findings are also consistent across repetitive reasoning-guided expansion experiments (Appendix B.4). Perhaps unsurprisingly, expanding based on consent provides the largest benefits. In addition, both datasets benefit from clarification on the privacy of information code; i.e., clarifying whether or not the attribute is sensitive. For PrivacyLens+, which follows the contextual integrity framework, we also observe the benefits of clarifying the alignment with norms code for information sharing, showing that we can improve the privacy judgment of LLMs by specifying the contextual norms of the information flows.
@@
Figure 6: The frequency with which each dataset field is selected when original scenarios are expanded based on each code using Gemini 2.5 Pro. Each code mainly selects one or two fields that are most relevant to the code definition to expand. Meanwhile, all code expansions are highly concentrated among a small subset of fields – transmission principle for PrivacyLens+, and detail and reveal reason for ConfAIde+, which indicates the sources of contextual ambiguity, pinpointing this small subset of fields as the primary origin.
@@
To discover where context ambiguity stems from, and understand what drives the performance gains observed through reasoning-guided expansion, we highlight in Figure 6 the frequency with which each field is selected for reasoning-guided expansion for both datasets. We observe that each code mainly focuses on one or two fields to expand, and a small number of fields that contribute to most of the context ambiguity – transmission principle for PrivacyLens+, and detail and reveal reason for ConfAIde+– are selected by the model to expand most codes. This shows that models are able to reason not only about the source of ambiguity, but also about which dataset fields are the most contextually ambiguous. Our reasoning-guided expansion framework also highlights transmission principle as the most ambiguous contextual integrity parameter in PrivacyLens+, providing meaningful context disambiguation directions that improve privacy assessments in the future.
@@
Many of the reasoning-guided codes we develop (see Table 2) apply to only a small subset of examples, but we observe in Figure 5 that expanding context based on a single code significantly improves performance across all examples. To understand how these codes generalize, we split the 120 coded examples in PrivacyLens+ into two groups, those labeled with a particular code and those without. Figure 7 shows the performance of those two groups before and after expansion using each of the 9 codes. Perhaps surprisingly, the performance gains (measured by the large percentage of right -> wrong group) on examples labeled with a particular code are not significantly higher than the performance gains on those without (i.e., clarification in one code generalizes in terms of helping in examples where the LLM-generated reasoning stated other assumptions). This shows that on a per-example basis, expanding based on the code mentioned in LLM-generated reasoning does not outperform choosing any other codes from the codebook. This is likely due to LLM-provided reasoning not reflecting the underlying reasoning that led to the specific judgment, as also reported in prior work (chen2025reasoning, ). Additionally, many labeled examples may have more than one code assigned to them, making it harder to dissect the effect of one particular code. Nevertheless, model-generated reasoning in aggregate reflects the underlying considerations in the LLM reasoning process, as supported by the observation that for all codes, the expansions correct most of the previously incorrectly judged examples (i.e., 27% - 70% examples are right -> wrong). This suggests that while aggregate reasoning provides a powerful signal for identifying general areas of ambiguity, the specific rationale provided for any single decision may not be a reliable guide to the model’s underlying inference process, highlighting a critical challenge in LLM interpretability.
@@ 7 Discussion and Limitations
It is unclear whether the LLM expansions of context are representative of real-world scenarios, due to both the synthetic generation of data and the instructions provided. Moreover, despite our best efforts, it’s unclear if the added context is even consistent with the scenario. The results do show, however, that context plays a key role in terms of substantially improving privacy decision-making and prompt sensitivity.
@@
The results in this paper are generated with approximately 200k LLM inference invocations with on the order of 100 input tokens and 10 output tokens per invocation. We required 35k Gemini 2.5 Pro invocations to create PrivacyLens+ and 23k invocations to create ConfAIde+. Our results required 137k calls to Gemini and 135k calls to both GPT and Claude. The total number of inference invocations as part of this project are likely an order of magnitude higher due to changes in experiment design, prompts, and other dead ends.

### v2

@@ 5.1 Enabling context disambiguation across datasets
Our context expansions are designed to simulate real-life clarifications a personal agent could obtain from a user, which can be label-independent (e.g., the data sender John is a mid-level manager at a startup) or label-dependent (e.g., the data subject Emily gave John her consent to share her information). In label-dependent expansions, we generate contexts designed to align with the example’s appropriateness label. Specifically, we generate inappropriate contexts (denoted $\mathcal{N^{-}}$) for $\mathcal{S^{-}}$ examples to reinforce their inappropriateness and appropriate contexts (denoted $\mathcal{N^{+}}$) for $\mathcal{S^{+}}$ examples to reinforce their appropriateness. Crucially, these label-dependent expansions, which deliberately reinforce the example’s initial appropriateness label, are distinct from user clarifications in authentic user-agent interactions. While real-word user clarifications typically aim to bolster agent decision confidence and can shift the perceived appropriateness in either direction, our expansions are not intended to mirror such interactions. Instead, they are specifically constructed to systematically evaluate how LLM privacy decision-making responds to additional contextual information generated via various strategies.
@@ 5.2 Reasoning-guided expansion
The reasoning provided by models alongside appropriateness judgments reveals that they frequently make assumptions when context is under-specified. Next, we aim to uncover these implicit assumptions and subsequently generate context expansions that obviate model reliance on them. As we show in Section 6, this enables models to become significantly more effective in their privacy judgments.
@@
Suitability of communication channel (Suitability) Whether sending the data over this specific communication channel is okay $49$ 0
@@
Alignment with norms (Norms) Whether it is a standard practice to share this data in this context $29$ 77
Purpose (Purpose) Whether there is an intended goal or specific reason for sharing or utilizing the information $16$ 77
Recipient authorization (Recipient Auth) Whether the intended recipient is permitted to access or receive the information $15$ 0
Established practices (Practices) Whether sharing this information follows a previous approach or agreement between the parties $11$ 44
Sender authorization (Sender Auth) Whether the sender is permitted to share or transmit the information 55 22
Safety guidelines (Safety) Whether this information violates the content moderation policies of the platform on which it’s being shared 33 0
@@
Building on extensive research into LLM reasoning geva2021aristotle; sprague2023musr; yuchen2025zebralogic; jacovi2024chain and factors influencing their outputs, we seek to understand appropriateness decisions of LLMs by prompting them to provide concise reasoning alongside each yes/no decision (our prompts can be found in Appendix E). Although elicited reasoning may not perfectly represent internal processes for models, it offers valuable insights into their decision-making chen2025reasoning. This reasoning elicitation step yielded performance comparable to that of models not providing explicit reasoning (P=83.5%, R=74.4%, and $F_{1}$=78.7% for PrivacyLens+, and P=97.6%, R=75.9%, and $F_{1}$=85.4% for ConfAIde+).
To understand model reasoning and identify implicit assumptions, two authors analyzed 40 outputs for PrivacyLens+ (10 per category: TPs, TNs, FPs, FNs) and 21 outputs for ConfAIde+ (7 each for TPs, TNs, FNs; FPs were excluded due to insufficient samples). They examined the prompts and reasoning for those examples—blind to ground-truth labels and LLM judgments—to identify assumptions and develop the nine codes shown in Table 2. Subsequently, using this established codebook, the same two authors independently coded a larger, stratified random sample of 120 PrivacyLens+ examples (30 per category) and 50 ConfAIde+ examples (15 TPs, TNs, FNs, and 5 FPs). Disagreement was resolved through discussion to achieve consensus cascio2019team. Table 2 presents the resulting code occurrences across datasets.
@@
To test if codes are more effective at reducing implicit reasoning assumptions, we prompt the LLM to generate label-dependent contexts based on each code (denoted as $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$) for all $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$ examples respectively in both datasets. Our instruction specifically asks the LLM to first identify the most suitable field to expand based on the code and its definition, and then to expand the selected field following this code to make the $\mathcal{S^{-}}$ more inappropriate, or $\mathcal{S^{+}}$ more appropriate. Similarly to Shao et al. salt-nlp2024privacylens, we explicitly ask the LLM to add descriptive contexts and avoid using evaluative words like “sensitive” and “non-sensitive” that would introduce reasoning shortcuts in the expansions. As we did for our label-dependent expansion of each field, for ConfAIde+ we again need to further integrate $\mathcal{N^{-}_{C}}$ and $\mathcal{N^{+}_{C}}$ into $\mathcal{S^{-}}$ and $\mathcal{S^{+}}$. In total, we generate an additional $9\times 493=$ 4,437 pairs of $\mathcal{S^{-}_{N^{-}_{C}}}$ and $\mathcal{S^{+}_{N^{+}_{C}}}$ examples for PrivacyLens+, and $9\times 270=$ 2,430 pairs for ConfAIde+. We provided detailed descriptions of our generation procedure in Appendix D.1.3 and  D.2.4.
In total, we generate 19,720 examples for PrivacyLens+, and 12,960 examples for ConfAIde+  [^2] Our dataset is available at https://github.com/google-parfait/contextual-privacy-and-security. To assess the quality of the expansions obtained through Camber, we manually analyzed a subset of the examples, discovering that for both datasets, 97% of expansions are plausible and in line with the original context, reasoning codes and appropriateness labels. The analysis process and detailed results are described in Appendix B.3.
@@ 6 Results
To measure the benefits of reasoning-guided contextual disambiguation we ask the following questions: (1) how reasoning-guided expansion performance compares to other expansion strategies, and (2) how well the reasoning-guided expansion codes generalize. We provide in-depth quantitative and qualitative performance assessments in Appendix B.
@@
Figure 4 compares the performance of privacy judgments on the reasoning-guided expansion with that on the original scenarios, and the label-independent and label-dependent expansions. For each expansion, we report the average $F_{1}$ across all fields/codes for that expansion and the $F_{1}$ for the top-performing field/code. Looking at the average expansion performance, we observe that the label-independent expansion provides little utility in improving performance over the no expansion baseline. In contrast, the label-dependent expansion shows consistent improvements across models, especially for PrivacyLens+. This is likely due to the expansion being based on contextual integrity: forcing context disambiguation to operate within the contextual integrity parameters provides clear benefits in improving appropriateness judgments. Nevertheless, the reasoning-guided expansion outperforms the others for both datasets, which further reinforces the benefits of basing context disambiguation on the codes that we introduced in Section 5.2.
@@
Figure 4 also highlights the performance of the top-performing field / code for each expansion, revealing a high degree of consistency in the fields / codes chosen by each model. On the PrivacyLens+ dataset, the agreement is unanimous, with all three models having identical top-performing fields / codes for each respective strategy. This pattern of consensus is also observed for ConfAIde+, where there is strong agreement for both the label-dependent and reasoning-guided expansions.
@@
Analyzing the performance of the top-performing fields and codes clearly highlights the utility of reasoning-guided expansion: it delivers substantial performance increase and reduction in prompt sensitivity that is also consistent across models, compared to the other expansion strategies. In addition, the expansion codes that we develop using PrivacyLens+ are also able to increase performance on the ConfAIde+ dataset, highlighting their generalization capability. A detailed breakdown and discussion of the performance across individual fields and codes is available in Appendices B.1 and B.2.
Entropy reduction for LLM judgments.
To quantitatively measure context ambiguity, we analyze the variance in the LLM’s privacy judgments by prompting Gemini 2.5 Pro 100 times for each example with temperature set to 1.0. We then calculate the Shannon entropy across the 100 privacy judgments for each example. Results are shown in Figure 5 for PrivacyLens+ and ConfAIde+ where higher entropy indicates higher variance in the responses, and hence greater ambiguity.
Our analysis of the original, no expansion examples reveals a notable disparity based on the appropriateness labels of the examples: the mean entropy of inappropriate class (0.089 for PrivacyLens and 0.014 for ConfAIde) is significantly lower compared to appropriate class (0.227 and 0.084, respectively). This indicates that the model exhibits higher consistency when denying data sharing when inappropriate than allowing sharing when appropriate. The finding complements the results in Table 1 which show higher precision across the board (85.6 - 95.6% for PrivacyLens, 98.6 - 100% for ConfAIde) than recall (40.6 - 75.7% and 33 - 78.5%, respectively). We also find that entropy is lower on privacy judgments for ConfAIde compared to PrivacyLens. This can be attributed to the scenario formats: the plain-text description in ConfAIde encodes more context and hence less ambiguity than the 5 contextual integrity parameters in PrivacyLens.
Figure 5: The entropy scores of the LLM judgments, for no expansion examples (No Exp) and reasoning-guided expansion examples across both datasets. For each example, the entropy score is computed from the probability distribution of 100 privacy judgments generated by Gemini 2.5 Pro with temperature 1.0.
We repeat this analysis on the reasoning-guided expansion examples generated by Camber based on the codes identified in  2. The results for PrivacyLens+ show that context disambiguation reduces entropy for both inappropriate and appropriate scenarios compared to their no expansion baselines. For ConfAIde+, however, the reduction in entropy following reasoning-guided expansion is more limited, likely due to the already low context ambiguity in original examples. Nevertheless, the expanded ConfAIde+ examples still achieve the lowest entropy scores overall, confirming the utility of the reasoning-guided strategy in reducing context ambiguity.
Entropy reduction for human judgments.
To validate whether the context ambiguity observed in LLM responses is also reflected in annotator disagreement, the six authors are tasked with labeling scenarios from the PrivacyLens+ dataset. We sample 180 pairs of no expansion and reasoning-guide expansion examples for each of the 18 code-label combinations (9 privacy codes $\times$ 2 appropriateness labels). To prevent annotator bias, the six authors are split into two groups of three. For each example pair, one group has to label the no expansion example, while the second group labels the code expansion counterpart. This ensures no annotator sees both examples belonging to the same pair. Each of the 360 examples is labeled three times. We then calculate entropy from the three annotations for each example to quantify disagreement.
The results, illustrated in Figure 6, are consistent with the LLM-based entropy analysis. The annotators exhibit higher agreement (lower entropy) when judging inappropriate examples compared to appropriate ones. Furthermore, the reasoning-guided expansions significantly reduce entropy, and thus ambiguity, for both classes. This reduction is observed across all 9 codes for the inappropriate examples and across 7 out of 9 codes for the appropriate ones. Notably, ambiguity is eliminated entirely (entropy = 0) in several cases, including appropriate examples under the consent code (e.g., "the data subject has consented") as well as inappropriate examples under the established practices (e.g., "no established practices for sharing") and recipient authorization (e.g., "the recipient is not authorized").
These findings validate our use of LLM entropy as a proxy for context ambiguity. They also demonstrate that the reasoning-guided expansion strategy is effective at disambiguating context for both LLMs and human evaluators.
Figure 6: The entropy scores of the human judgments, for no expansion examples and reasoning-guided expansion examples across PrivacyLens+. For each example, the entropy score is computed from three independent human judgments.
@@
Figure 7: The frequency with which each dataset field is selected when original scenarios are expanded based on each code using Gemini 2.5 Pro. Each code mainly selects one or two fields that are most relevant to the code definition to expand. Meanwhile, all code expansions are highly concentrated among a small subset of fields – transmission principle for PrivacyLens+, and detail and reveal reason for ConfAIde+, which indicates the sources of context ambiguity, pinpointing this small subset of fields as the primary origin.
@@
To discover where context ambiguity stems from and understand what drives the performance gains observed through reasoning-guided expansion, we highlight in Figure 7 the frequency with which each field is selected for reasoning-guided expansion for both datasets. We observe that each code mainly focuses on one or two fields to expand, and a small number of fields that contribute to most of the context ambiguity – transmission principle for PrivacyLens+, and detail and reveal reason for ConfAIde+– are selected by the model to expand most codes. This shows that models are able to reason not only about the source of ambiguity, but also about which dataset fields are the most contextually ambiguous. Our reasoning-guided expansion framework also highlights transmission principle as the most ambiguous contextual integrity parameter in PrivacyLens+, providing meaningful context disambiguation directions that improve privacy assessments in the future.
@@
Many of the reasoning-guided codes we develop (see Table 2) apply to only a small subset of examples, but we observe in Figure 4 that expanding context based on a single code significantly improves performance across all examples. To understand how these codes generalize, we split the 120 coded examples in PrivacyLens+ into two groups, those labeled with a particular code and those without. Figure 8 shows the performance of those two groups before and after expansion using each of the 9 codes. Perhaps surprisingly, the performance gains (measured by the large percentage of wrong -> right group) on examples labeled with a particular code are not significantly higher than the performance gains on those without (i.e., clarification in one code generalizes in terms of helping in examples where the LLM-generated reasoning stated other assumptions). This shows that on a per-example basis, expanding based on the code mentioned in LLM-generated reasoning does not outperform choosing any other codes from the codebook. This is likely due to LLM-provided reasoning not reflecting the underlying reasoning that led to the specific judgment, as also reported in prior work (chen2025reasoning). Additionally, many labeled examples may have more than one code assigned to them, making it harder to dissect the effect of one particular code. Nevertheless, model-generated reasoning in aggregate reflects the underlying considerations in the LLM reasoning process, as supported by the observation that for all codes, the expansions correct most of the previously incorrectly judged examples (i.e., 27% - 70% examples are wrong -> right). This suggests that while aggregate reasoning provides a powerful signal for identifying general areas of ambiguity, the specific rationale provided for any single decision may not be a reliable guide to the model’s underlying inference process, highlighting a critical challenge in LLM interpretability. A qualitative analysis of the expansions is performed in Appendix B.3.
@@ 7 Discussion and Limitations
It is unclear whether the LLM expansions of context are representative of real-world scenarios, due to both the synthetic generation of data and the instructions provided. Moreover, our framework does not provide a solution for obtaining user clarifications in practice. Any production use of our framework must carefully consider potential harms, including the amplification of bias and fairness issues, and any such system should be developed using an interdisciplinary approach with expert feedback. Our results do show, however, that context plays a key role in terms of substantially improving privacy decision-making and reducing prompt sensitivity of agents.
@@
The results in this paper are generated with approximately 200k LLM inference invocations, using on the order of 100 input tokens and 10 output tokens per invocation. We required 35k Gemini 2.5 Pro invocations to create PrivacyLens+ and 23k invocations to create ConfAIde+. Our results required 137k calls to Gemini and 135k calls to both GPT and Claude. The total number of inference invocations as part of this project is likely an order of magnitude higher due to changes in experiment design, prompts, and other dead ends.